A produce distributor uses 789 packing crates a month, which it purchases at a cost of $12 each. The manager has assigned an annual carrying cost of 38 percent of the purchase price per crate. Ordering costs are $30. Currently the manager orders once a month.

How much could the firm save annually in ordering and carrying costs by using the EOQ? (Round intermediate calculations and final answer to 2 decimal places. Omit the "$" sign in your response.) Savings $ per year _______
Answer #1

Annual demand, D = 789 x 12 = 9,468 crates
Carryng cost, h = $12 x 38% = $4.56 per annum
Ordering cost, K = $30

So, EOQ = (2.D.K / h)1/2 = sqrt(2*9468*30 / 4.56) = 352.96 or 353 units

Total relevant cost under EOQ = (D/EOQ)*K + (EOQ/2)*h = (9468/353)*30 + (353/2)*4.56 = $1609.49

For monthly ordering policy, average order size = D/12 = 789 units

Total relevant cost = 12*30 + (789/2)*4.56 = $2,158.92

So, savings = $2,158.92 - $1609.49 = $549.43